An aluminum cup contains 225 g of water and a 40-g copper stirrer, all at 27°C. A 436-g sample of silver at an initial temperature of 86°C is placed in the water. The stirrer is used to stir the mixture gently until it reaches its final equilibrium temperature of 32°C. Calculate the mass of the aluminum cup.
____g
Answer #1

Let mAl be the mass of the aluminium cup.

Then,

[mAlCAl + mCuCCu + mwCw](32 - 27) = mAgCAg(86 - 32)

=> [mAl * 0.9 + 0.04 * 0.385 + 0.225 * 4.184] * 5 = 0.436 * 0.240 * 54

=> mAl = 0.193 kg = 193 g

[Reference for specific heats: http://www2.ucdsb.on.ca/tiss/stretton/database/Specific_Heat_Capacity_Table.html]